OpenStudy (kkutie7):

A commuter train travels between two stations by accelerating at the rate a for a time t1 and then slowing down with a negative acceleration of -5a for a time t2. The two stations are a distance d apart and the train never reaches its maximum speed. The train stops at each station. Find times t1 and t2.

OpenStudy (kkutie7):

I'm also not allowed to use any other equation except these: \[v^{\rightarrow}=\frac{dr^{\rightarrow}}{dt}\] \[a^{\rightarrow}=\frac{dv^{\rightarrow}}{dt}\] \[v^{\rightarrow}=v _{0}^{\rightarrow}+at^{\rightarrow}\] \[r^{\rightarrow}=r _{0}^{\rightarrow}+v _{0}^{\rightarrow}t+\frac{1}{2}a^{\rightarrow}t^{2}\]